OpenStudy (anonymous):

What is the slope-intercept form of the function that contains the point (8, 12) and has a slope of −2?

OpenStudy (anonymous):

@perl

OpenStudy (perl):

we can use the 'point slope' form for equation of line y-y1 = m(x - x1)

OpenStudy (perl):

there is more than one form for the equation of a line. Another form that comes up often is 'slope intercept' form, y = mx + b

OpenStudy (anonymous):

i can do the y = mx + b

OpenStudy (perl):

ok, then we will have to do some algebra

OpenStudy (perl):

we are given the slope is -2, so lets plug that in for m

OpenStudy (anonymous):

never done algebra before D:

OpenStudy (perl):

this is an algebra class , right?

OpenStudy (anonymous):

math

OpenStudy (perl):

what do you get when you plug in m = -2 we get y = -2x + b

OpenStudy (perl):

we still need to find 'b'

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

plz asap